OpenStudy (purexyz):

Rewrite the equation of a line in point-slope form to slope-intercept form: y-3=1/2(x-8)

OpenStudy (anonymous):

y - 3 = 1/2(x - 8) -- distribute through the parenthesis y - 3 = 1/2x - 4 -- add 3 to both sides y = 1/2x - 4 + 3 y = 1/2x - 1 <-- slope intercept form

OpenStudy (purexyz):

thanks! : )